# Hypothesis testing for community structure in temporal networks using e-values

### v2

@@ 4.3.1 Model
While the previous model allows for correlation between successive realizations of the network, it does not allow the node communities to change with time. To allow for this property, we discuss the dynamic SBM (Matias and Miele,, 2017). Recall that we represent our temporal network as a sequence of adjacency matrices ${\bf A}^{(1)},\dots,{\bf A}^{(T)}$. Additionally, we have community labels $\boldsymbol{c}^{(1)},\dots,\boldsymbol{c}^{(T)}$ such that $\boldsymbol{c}^{(t)}\in\{1,2,\dots,K\}^{n}$ and $c_{i}^{(t)}=k$ means that node $i$ is in group $k$ at time $t$. Thus, the community labels are now allowed to vary with time. We also define $\boldsymbol{c}_{i}=(c_{i}^{(1)},\dots,c_{i}^{(T)})^{\top}$ to represent the group membership for node $i$ over time.
Matias and Miele, (2017) assume that $\boldsymbol{c}_{1},\dots,\boldsymbol{c}_{n}$ are independent and identically distributed random variables. For a given node, however, $\boldsymbol{c}_{i}$ is a Markov chain defined by transition matrix $\boldsymbol{\pi}\in(0,1)^{K\times K}$ and initial stationary distribution $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{K})^{\top}$. In words, if node $i$ is in community $k$ at time $t$, then the probability that it is in group $k^{\prime}$ at time $t+1$ is $\pi_{k,k+1}$ for $k,k^{\prime}\in\{1,\dots,K\}$. Given $\boldsymbol{c}^{(1)},\dots,\boldsymbol{c}^{(T)}$, we generate a SBM as before:
@@ 4.4.1 Model
For the final simulation, we consider a temporal degree-corrected block model (DCBM). The DCBM was first proposed in Karrer and Newman, (2011) to model the fact that many real-world networks have degree heterogeneity. As before, let $\boldsymbol{c}$ correspond to the community labels and $\bf B$} represent the community edge probabilities. We also introduce node weights, $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})^{\top}$ such that if $\theta_{i}$ is large, then we expect node $i$ to have relatively more edges. Then the probability distribution of the network is defined as
@@
Thus, the probability of an edge between nodes $i$ and $j$ depends not only on their community assignment (through $\bf B$}) but also their relative importance (via $\theta$). Note that if $\theta_{i}\equiv\theta$ for all $i$, then we recover the non-degree-corrected SBM. Furthermore, if $B_{k,k^{\prime}}=b$ for all $k,k^{\prime}$, then the DCBM reduces to the Chung-Lu (CL) model (Chung and Lu,, 2002) which does not possess community structure, but also differs from the ER model, i.e., $P_{ij}\propto\theta_{i}\theta_{j}$.
There are various ways for this model to vary temporally including through the community assignments $(\boldsymbol{c})$, block probabilities $({\bf B})$ or weight parameters $(\boldsymbol{\theta})$ (Wilson et al.,, 2019). For simplicity, we follow the approach of Matias and Miele, (2017) to let the community labels change with time, in addition to varying $\theta_{i}$ (Wilson et al.,, 2019). Specifically, we generate
@@ 4.4.3 Results
The results are similar to the previous ones. As $\delta$ and $n$ increase, the median e-value also increases. When $\delta$ is varied, the max and $\kappa=0.25$ calibrators yield the largest e-values for all values of $\delta$. Again, the $\kappa=0.75$ calibrator tends to have the smallest e-value. The main difference from the previous settings is that the e-value increase occurs for smaller $\delta$ and $n$. This is because the added heterogeneity from the $\boldsymbol{\theta}$ parameters means the model diverges from the ER model more quickly than a basic SBM. Indeed, in Figure 3(a), the median e-values are greater than 1 even when $\delta=0$ (no block model structure). Even though $\delta=0$ and the null hypothesis appears to be true, the DCBM behaves like a CL model which still differs from the ER model, causing the slight inflation in the e-value. This is a key limitation of the method from Bickel and Sarkar, (2016) in that it only tests against the ER null. As seen here, it is possible for a network to diverge from the ER null without possessing a community structure. This idea will be more fully explored in the real-data analysis.
@@ 5.1 Choice of null model
We turn our attention to studying real-world networks, further demonstrating the ability of the proposed test to combine information across dependent network realizations. In the previous sections, we leveraged the static community structure hypothesis test from Bickel and Sarkar, (2016). The proposed e-value testing procedure, however, works more generally for any test that yields p-values on static networks. Indeed, in many cases, the ER model may not be a satisfactory null model as it is too easy to reject. In other words, almost all real-world networks diverge from a simple ER model, even if they do not have a community structure. Specifically, in the dynamic DCBM simulations, we found that even if there was no block community structure $(\delta=0)$, the CL model alone led to slightly larger e-values then we would have expected under the null (ER model). Moreover, in the following examples, we found that each network (save Kenya) yielded an infinite e-value when using the method from Bickel and Sarkar, (2016). If every network yields an infinite e-value, then it may not be useful metric for discriminating community strength in networks.
In light of these observations, we apply the bootstrap testing approach from Yanchenko and Sengupta, (2024) in order to leverage the Chung-Lu (CL) model as the null model. We prefer this null as it can model degree heterogeneity while still not exhibiting community structure, making it a more flexible null model. Indeed, the CL model is a close cousin of the configuration model which is commonly used as a null model in community detection algorithms (e.g., Newman,, 2006).
@@ 5.2 Bootstrap method
We briefly describe the bootstrap testing method from Yanchenko and Sengupta, (2024). The authors begin by defining the expected edge density difference (E2D2) parameter to quantify the strength of a community structure. Specifically, for adjacency matrix $\bf A$} and community labels $\boldsymbol{c}$, the E2D2 statistic is defined as scaled difference between the intra- and inter-community edge density, i.e.,
@@
The bootstrap test proceeds as follows. Given the observed adjacency matrix $\bf A$}, $\tilde{U}({\bf A})$ is calculated as in (7). Since we are testing against the null hypothesis of a CL model, we need to generate CL networks that resemble $\bf A$}. To do this, we first estimate the weight vector $\boldsymbol{\theta}$ with using the rank 1 adjacency spectral embedding (ASE) (Sussman et al.,, 2012):
$\boldsymbol{\theta}=|\lambda_{1}|^{1/2}\,{\bf u}$
where $\lambda_{1}$ is the largest eigenvalue of $\bf A$} (in magnitude) and $\bf u$} is the corresponding eigenvector. Then for the $b$th bootstrap iteration, we sample from $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})^{\top}$ with replacement to obtain $\boldsymbol{\theta}^{(b)}=(\theta_{1}^{(b)},\dots,\theta_{n}^{(b)})^{\top}$ and then sample a network ${\bf A}^{(b)}$ using $\boldsymbol{\theta}^{(b)}$. We compute $\tilde{U}({\bf A}^{(b)})=\max_{\boldsymbol{c}}\{U({\bf A}^{(b)},\boldsymbol{c})\}$ and repeat this process for $B$ iterations. Finally, the p-value is
@@
See Yanchenko and Sengupta, (2024) for further details and properties of this test. We briefly note that only the largest connected component of the graph is used when computing $\tilde{U}({\bf A})$ and the number of communities $K$ is found using the Fast Greedy Algorithm, with an upper-bound set at $\tilde{n}^{1/2}$, where $\tilde{n}$ is the number of nodes in the largest connected component. In this real-data analysis, we use this test to compute the p-value on each network snapshot, and then flip to e-values using the procedure described in Section 3.1.
@@ 5.3 Data sets
To study the performance of the proposed hypothesis test under various settings, we consider five real-world networks of various sizes and domains. The first four networks are all proximity networks which record an edge if two people are close to each other: Kenya (Kiti et al.,, 2016), Reality (Eagle and Pentland,, 2006), Hospital (Vanhems et al.,, 2013) and High School 1 (Mastrandrea et al.,, 2015). The remaining network is constructed from online communication on a social media platform: College (Panzarasa et al.,, 2009). For each network, we apply the bootstrap hypothesis test with CL null from Yanchenko and Sengupta, (2024) with $B=1,000$ bootstrap iterations. We also look at the effect of the number of snapshots by varying $T\in\{5,10\}$, where we again assume that this choice was made before the data were collected. We emphasize that the temporal network is given with each edge having a time stamp. Thus, the total observation time of the network process is fixed, but we can vary the number of snapshots based on how we bin the data. Indeed, increasing $T$ should not be confused with increasing the length of the observation period in this setting.
@@ 5.4 Results
For each network, the number of nodes $n$ and e-value are reported in Table 1. We use the $\kappa=0.25$ calibrator because it yielded the largest e-values in the simulation studies when $E>1$. Kenya does not show evidence of a statistically significant community structure given the small e-values. Hospital and College have slightly larger e-values, but still not large enough to warrant rejection of the null hypothesis. For Hospital, this finding accords with that in Yanchenko and Sengupta, (2024), which found no community structure compared to the CL null for this network, though this was when the network was treated as static, i.e., $T=1$. For Reality and High School 1, however, there is very strong evidence to reject the null hypothesis and claim that these networks exhibit a statistically meaningful community structure. In general, the results do not seem to be overly sensitive to the choice of $T$. Regardless, we stress that in practice this choice should be carried out in a meaningful way based on the domain.
@@ 6 Conclusion
This raises an important question which we have hitherto not explicitly addressed: what does it mean for a temporal network to have community structure? This question is discussed in Cazabet and Rossetti, (2023). Perhaps the simplest definition is fixed community structure where the labels and underlying strength of the community structure are invariant with time. The correlated SBM discussed in Section 4.2 would be one example of this. Next, we may have persistent community structure where the community labels are now allowed to change over time, but the strength of the community structure is still roughly constant. The dynamic SBM (Section 4.3) and dynamic DCBM (Section 4.4) both exhibit this kind of community structure, as did the High School 1 network. The proposed hypothesis testing formulation is sensible for either of these cases.
@@
Another related possibility is accumulating community structure where the strength of the community structure increases with time. An example of this case comes from models like that of Sasahara et al., (2021). In this work, the authors propose a model for echo chambers formation in online social networks, describing how starting from a relatively well-mixed population can quickly lead to strong communities if edges are formed and removed based on some simple heuristics. In this case, the community structure is, by design, not constant, and only arises after a certain point. In this case, we may be more interested in change point detection to know when a community structure arises (e.g., Wilson et al.,, 2019).
@@
Regardless, a precise characterization of temporal network community structure is needed to develop a sensible hypothesis test. For example, is the intermittent community structure like that of Reality a meaningful structure? How “intermittent” can it be until it is no longer community structure? Does it make sense to test for community structure where communities are coming, going, merging and splitting through the process? In this work, we proposed a statistically valid test by extending the definition from static networks, but a more nuanced definition may be needed depending on the particular behavior of the network of interest. These and similar questions will also need to be addressed if the methods in this work are extended to test for other network properties, e.g., core-periphery structure (Yanchenko and Sengupta,, 2023).
@@
As another possible direction for future work, our proposed framework can easily be extended to other situations where multiple realizations of the network arise, i.e., multilayer networks. (Kivelä et al.,, 2014). For example, there could be a network where nodes are researchers, and one layer of the network represents collaborations as edges, a different layer uses edges to denote citations, and a third layer corresponds to a friendship network. If the practitioner is interested in testing for the significance of community structure across these different layers, then our approach could easily be applied.
@@
On the theoretical and methodological side, one possible improvement would be to construct the network-specific e-values directly, rather than applying a calibrator to a known p-value. The intuition would go roughly as follows. The individual p-values are themselves relatively efficient, and the “dream-world” e-value would be simply the reciprocal of that efficient p-value. Unfortunately, the reciprocal of a p-value is not an e-value, and, roughly speaking, the calibrator strategically inflates the p-value by a small amount before taking the reciprocal and returning an e-value. It is this preliminary inflation that, while validity is preserved, leads to a less-efficient-than-necessary e-value. A direct e-value construction like in Grünwald et al., (2024); Larsson et al., (2025) would generally be more efficient, and the construction of such an e-value is the focus of future research.
@@
We describe a hypothesis test for community structure on static networks from Bickel and Sarkar, (2016). Let $\mathcal{G}\equiv\mathcal{G}^{(1)}$ be a static network and ${\bf A}\equiv{\bf A}^{(1)}$ the associated adjacency matrix and define the Erdos-Renyi model as the null model. Specifically, define $\hat{p}$ as the estimated probability of an edge between any two nodes and let
@@
We include additional simulation results for the correlated SBM, dynamic SBM and dynamic DCBM in Figures 5, 6, 7, respectively. The trends are similar to those in the main manuscript.

### v3

@@ 4.3.1 Model
While the previous model allows for correlation between successive realizations of the network, it does not allow the node communities to change with time. To allow for this property, we discuss the dynamic SBM (Matias and Miele, 2017). Recall that we represent our temporal network as a sequence of adjacency matrices ${\bf A}^{(1)},\dots,{\bf A}^{(T)}$. Additionally, we have community labels $\boldsymbol{c}^{(1)},\dots,\boldsymbol{c}^{(T)}$ such that $\boldsymbol{c}^{(t)}\in\{1,2,\dots,K\}^{n}$ and $c_{i}^{(t)}=k$ means that node $i$ is in group $k$ at time $t$. Thus, the community labels are now allowed to vary with time. We also define $\boldsymbol{c}_{i}=(c_{i}^{(1)},\dots,c_{i}^{(T)})^{\top}$ to represent the group membership for node $i$ over time.
Matias and Miele (2017) assume that $\boldsymbol{c}_{1},\dots,\boldsymbol{c}_{n}$ are independent and identically distributed random variables. For a given node, however, $\boldsymbol{c}_{i}$ is a Markov chain defined by transition matrix $\boldsymbol{\pi}\in(0,1)^{K\times K}$ and initial stationary distribution $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{K})^{\top}$. In words, if node $i$ is in community $k$ at time $t$, then the probability that it is in group $k^{\prime}$ at time $t+1$ is $\pi_{k,k+1}$ for $k,k^{\prime}\in\{1,\dots,K\}$. Given $\boldsymbol{c}^{(1)},\dots,\boldsymbol{c}^{(T)}$, we generate a SBM as before:
@@ 4.4.1 Model
For the final simulation, we consider a temporal degree-corrected block model (DCBM). The DCBM was first proposed in Karrer and Newman (2011) to model the fact that many real-world networks have degree heterogeneity. As before, let $\boldsymbol{c}$ correspond to the community labels and $\bf B$} represent the community edge probabilities. We also introduce node weights, $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})^{\top}$ such that if $\theta_{i}$ is large, then we expect node $i$ to have relatively more edges. Then the probability distribution of the network is defined as
@@
Thus, the probability of an edge between nodes $i$ and $j$ depends not only on their community assignment (through $\bf B$}) but also their relative importance (via $\theta$). Note that if $\theta_{i}\equiv\theta$ for all $i$, then we recover the non-degree-corrected SBM. Furthermore, if $B_{k,k^{\prime}}=b$ for all $k,k^{\prime}$, then the DCBM reduces to the Chung-Lu (CL) model (Chung and Lu, 2002) which does not possess community structure, but also differs from the ER model, i.e., $P_{ij}\propto\theta_{i}\theta_{j}$.
There are various ways for this model to vary temporally including through the community assignments $(\boldsymbol{c})$, block probabilities $({\bf B})$ or weight parameters $(\boldsymbol{\theta})$ (Wilson et al., 2019). For simplicity, we follow the approach of Matias and Miele (2017) to let the community labels change with time, in addition to varying $\theta_{i}$ (Wilson et al., 2019). Specifically, we generate
@@ 4.4.3 Results
The results are similar to the previous ones. As $\delta$ and $n$ increase, the median e-value also increases. When $\delta$ is varied, the max and $\kappa=0.25$ calibrators yield the largest e-values for all values of $\delta$. Again, the $\kappa=0.75$ calibrator tends to have the smallest e-value. The main difference from the previous settings is that the e-value increase occurs for smaller $\delta$ and $n$. This is because the added heterogeneity from the $\boldsymbol{\theta}$ parameters means the model diverges from the ER model more quickly than a basic SBM. Indeed, in Figure 3(a), the median e-values are greater than 1 even when $\delta=0$ (no block model structure). Even though $\delta=0$ and the null hypothesis appears to be true, the DCBM behaves like a CL model which still differs from the ER model, causing the slight inflation in the e-value. This is a key limitation of the method from Bickel and Sarkar (2016) in that it only tests against the ER null. As seen here, it is possible for a network to diverge from the ER null without possessing a community structure. This idea will be more fully explored in the real-data analysis.
@@ 5.1 Choice of null model
We turn our attention to studying real-world networks, further demonstrating the ability of the proposed test to combine information across dependent network realizations. In the previous sections, we leveraged the static community structure hypothesis test from Bickel and Sarkar (2016). The proposed e-value testing procedure, however, works more generally for any test that yields p-values on static networks. Indeed, in many cases, the ER model may not be a satisfactory null model as it is too easy to reject. In other words, almost all real-world networks diverge from a simple ER model, even if they do not have a community structure. Specifically, in the dynamic DCBM simulations, we found that even if there was no block community structure $(\delta=0)$, the CL model alone led to slightly larger e-values then we would have expected under the null (ER model). Moreover, in the following examples, we found that each network (save Kenya) yielded an infinite e-value when using the method from Bickel and Sarkar (2016). If every network yields an infinite e-value, then it may not be useful metric for discriminating community strength in networks.
In light of these observations, we apply the bootstrap testing approach from Yanchenko and Sengupta (2024) in order to leverage the Chung-Lu (CL) model as the null model, i.e., testing $H_{0}:\text{Chung-Lu model}$ versus $H_{1}:\text{not }H_{0}$. We prefer this null as it can model degree heterogeneity while still not exhibiting community structure, making it a more flexible null model. Indeed, the CL model is a close cousin of the configuration model which is commonly used as a null model in community detection algorithms (e.g., Newman, 2006).
@@ 5.2 Bootstrap method
We briefly describe the bootstrap testing method from Yanchenko and Sengupta (2024). The authors begin by defining the expected edge density difference (E2D2) parameter to quantify the strength of a community structure. Specifically, for adjacency matrix $\bf A$} and community labels $\boldsymbol{c}$, the E2D2 statistic is defined as scaled difference between the intra- and inter-community edge density, i.e.,
@@
The bootstrap test proceeds as follows. Given the observed adjacency matrix $\bf A$}, $\tilde{U}({\bf A})$ is calculated as in (7). Since we are testing against the null hypothesis of a CL model, we need to generate CL networks that resemble $\bf A$}. To do this, we first estimate the weight vector $\boldsymbol{\theta}$ with using the rank 1 adjacency spectral embedding (ASE) (Sussman et al., 2012):
$\hat{\boldsymbol{\theta}}=|\lambda_{1}|^{1/2}\,{\bf u}$
where $\lambda_{1}$ is the largest eigenvalue of $\bf A$} (in magnitude) and $\bf u$} is the corresponding eigenvector. Then for the $b$th bootstrap iteration, we sample from $\hat{\boldsymbol{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{n})^{\top}$ with replacement to obtain $\boldsymbol{\theta}^{(b)}=(\theta_{1}^{(b)},\dots,\theta_{n}^{(b)})^{\top}$ and then sample a network ${\bf A}^{(b)}$ using $\boldsymbol{\theta}^{(b)}$. We compute $\tilde{U}({\bf A}^{(b)})=\max_{\boldsymbol{c}}\{U({\bf A}^{(b)},\boldsymbol{c})\}$ and repeat this process for $B$ iterations. Finally, the p-value is
@@
See Yanchenko and Sengupta (2024) for further details and properties of this test. We briefly note that only the largest connected component of the graph is used when computing $\tilde{U}({\bf A})$ and the number of communities $K$ is found using the Fast Greedy Algorithm, with an upper-bound set at $\tilde{n}^{1/2}$, where $\tilde{n}$ is the number of nodes in the largest connected component. In this real-data analysis, we use this test to compute the p-value on each network snapshot, and then flip to e-values using the procedure described in Section 3.1.
@@ 5.3 Data sets
To study the performance of the proposed hypothesis test under various settings, we consider five real-world networks of various sizes and domains. The first four networks are all proximity networks which record an edge if two people are close to each other: Kenya (Kiti et al., 2016), Reality (Eagle and Pentland, 2006), Hospital (Vanhems et al., 2013) and High School 1 (Mastrandrea et al., 2015). The remaining network is constructed from online communication on a social media platform: College (Panzarasa et al., 2009). For each network, we apply the bootstrap hypothesis test with CL null from Yanchenko and Sengupta (2024) with $B=1,000$ bootstrap iterations. We also look at the effect of the number of snapshots by varying $T\in\{5,10\}$, where we again assume that this choice was made before the data were collected. We emphasize that the temporal network is given with each edge having a time stamp. Thus, the total observation time of the network process is fixed, but we can vary the number of snapshots based on how we bin the data. Indeed, increasing $T$ should not be confused with increasing the length of the observation period in this setting.
@@ 5.4 Results
For each network, the number of nodes $n$ and e-value are reported in Table 1. We use the $\kappa=0.25$ calibrator because it yielded the largest e-values in the simulation studies when $E>1$, and a rejection threshold of $\bar{E}_{T}>20$. Kenya does not show evidence of a statistically significant community structure given the small e-values. Hospital and College have slightly larger e-values, but still not large enough to warrant rejection of the null hypothesis. For Hospital, this finding accords with that in Yanchenko and Sengupta (2024), which found no community structure compared to the CL null for this network, though this was when the network was treated as static, i.e., $T=1$. For Reality and High School 1, however, there is very strong evidence to reject the null hypothesis and claim that these networks exhibit a statistically meaningful community structure. In general, the results do not seem to be overly sensitive to the choice of $T$. Regardless, we stress that in practice this choice should be carried out in a meaningful way based on the domain.
@@ 6 Conclusion
This raises an important question which we have hitherto not explicitly addressed: what does it mean for a temporal network to have community structure? This question is discussed in Cazabet and Rossetti (2023). Perhaps the simplest definition is fixed community structure where the labels and underlying strength of the community structure are invariant with time. The correlated SBM discussed in Section 4.2 would be one example of this. Next, we may have persistent community structure where the community labels are now allowed to change over time, but the strength of the community structure is still roughly constant. The dynamic SBM (Section 4.3) and dynamic DCBM (Section 4.4) both exhibit this kind of community structure, as did the High School 1 network. The proposed hypothesis testing formulation is sensible for either of these cases.
@@
Another related possibility is accumulating community structure where the strength of the community structure increases with time. An example of this case comes from models like that of Sasahara et al. (2021). In this work, the authors propose a model for echo chambers formation in online social networks, describing how starting from a relatively well-mixed population can quickly lead to strong communities if edges are formed and removed based on some simple heuristics. In this case, the community structure is, by design, not constant, and only arises after a certain point. In this case, we may be more interested in change point detection to know when a community structure arises (e.g., Wilson et al., 2019).
@@
Regardless, a precise characterization of temporal network community structure is needed to develop a sensible hypothesis test. For example, is the intermittent community structure like that of Reality a meaningful structure? How “intermittent” can it be until it is no longer community structure? Does it make sense to test for community structure where communities are coming, going, merging and splitting through the process? In this work, we proposed a statistically valid test by extending the definition from static networks, but a more nuanced definition may be needed depending on the particular behavior of the network of interest. These and similar questions will also need to be addressed if the methods in this work are extended to test for other network properties, e.g., core-periphery structure (Yanchenko and Sengupta, 2023).
@@
As another possible direction for future work, our proposed framework can easily be extended to other situations where multiple realizations of the network arise, i.e., multilayer networks. (Kivelä et al., 2014). For example, there could be a network where nodes are researchers, and one layer of the network represents collaborations as edges, a different layer uses edges to denote citations, and a third layer corresponds to a friendship network. If the practitioner is interested in testing for the significance of community structure across these different layers, then our approach could easily be applied.
@@
On the theoretical and methodological side, one possible improvement would be to construct the network-specific e-values directly, rather than applying a calibrator to a known p-value. The intuition would go roughly as follows. The individual p-values are themselves relatively efficient, and the “dream-world” e-value would be simply the reciprocal of that efficient p-value. Unfortunately, the reciprocal of a p-value is not an e-value, and, roughly speaking, the calibrator strategically inflates the p-value by a small amount before taking the reciprocal and returning an e-value. It is this preliminary inflation that, while validity is preserved, leads to a less-efficient-than-necessary e-value. A direct e-value construction like in Grünwald et al. (2024); Larsson et al. (2025) would generally be more efficient, and the construction of such an e-value is the focus of future research.
@@
We describe a hypothesis test for community structure on static networks from Bickel and Sarkar (2016). Let $\mathcal{G}\equiv\mathcal{G}^{(1)}$ be a static network and ${\bf A}\equiv{\bf A}^{(1)}$ the associated adjacency matrix and define the Erdos-Renyi model as the null model. Specifically, define $\hat{p}$ as the estimated probability of an edge between any two nodes and let
@@
Simulation study type I error and power results
In the main manuscript, we are primarily interested in reporting the raw e-value as a continuous measure of evidence for/against the null hypothesis. If a typical rejection decision is required, however, a binary threshold can be used. Towards this end, we report the rejection rate for the settings in the main manuscript using $\bar{E}_{T}>20$ as the rejection decision, corresponding to a type I error of no more than $\alpha=0.05$. The results are in Figures 4,5,6. When $\delta=0$, the null hypothesis is true and the test maintains a low rejection rate. For $\delta>0$, however, the alternative hypothesis is true and the test yields a large rejection rate.
(a) Varying $\delta$
(b) Varying $n$
Figure 4: Simulation rejection rates for correlated SBM networks with $\rho=0.25$ and rejection threshold $\bar{E}_{T}>20$. Dashed black line corresponds to $\alpha=0.05$.
(a) Varying $\delta$
(b) Varying $n$
Figure 5: Simulation rejection rates for dynamic SBM networks with $\boldsymbol{\pi}_{1}$ and rejection threshold $\bar{E}_{T}>20$. Dashed black line corresponds to $\alpha=0.05$.
(a) Varying $\delta$
(b) Varying $n$
Figure 6: Simulation rejection rates for dynamic DCBM networks with $\epsilon=0.6$ and rejection threshold $\bar{E}_{T}>20$. Dashed black line corresponds to $\alpha=0.05$.
@@
We include additional simulation results for the correlated SBM, dynamic SBM and dynamic DCBM in Figures 8, 9, 10, respectively. The trends are similar to those in the main manuscript.